### v1

# Towards automated formal security analysis of SAML V2.0 Web Browser SSO standard — the POST/Artifact use case

Abstract
Single Sign-On (SSO) protocols streamline user authentication with a unified login for multiple online services, improving usability and security. One of the most common SSO protocol frameworks — the Security Assertion Markup Language V2.0 (SAML) Web SSO Profile — has been in use for more than two decades, primarily in government, education and enterprise environments. Despite its mission-critical nature, only certain deployments and configurations of the Web SSO Profile have been formally analyzed. This paper attempts to bridge this gap by perforing a comprehensive formal security analysis of the SAML V2.0 SP-initiated SSO with POST/Artifact Bindings use case. Rather than focusing on a specific deployment and configuration, we closely follow the specification with the goal of capturing many different deployments allowed by the standard. Modeling and analysis is performed using Tamarin prover — state-of-the-art tool for automated verification of security protocols in the symbolic model of cryptography. Technically, we build a meta-model of the use case that we instantiate to eight different protocol variants. Using the Tamarin prover, we formally verify a number of critical security properties for those protocol variants, while identifying certain drawbacks and potential vulnerabilities.
Index Terms:
security protocols, formal proof, formal verification, SAML, security analysis, Single Sing-On, SSO, Tamarin prover
[…]
=-21pt
I Introduction
Increasing number of online services are requiring user authentication, with passwords still being the prevalent authentication method. Accessibility issues related to unique strong passwords naturally lead to password reuse, which, in turn, makes services vulnerable to various attacks [1, 2], degrading the overall security on the Internet. Simultaneously, the burden on service providers to deploy, maintain and safeguard the authentication infrastructure further compounds the complexity. In response to these challenges, the concept of Web Single Sign-On (SSO) emerged at start of this millennium. SSO protocols enable exchange of authentication and authorization data between trusted identity providers (IdPs) and service providers (SPs), enabling users to authenticate once at an IdP and access services across multiple SPs.
The Security Assertion Markup Language (SAML) version 2.0 was adopted as a standard in 2005 by the Organization for the Advancement of Structured Information Standards (OASIS), replacing version 1.0 adopted in 2002. Since that time, SAML V2.0 has emerged as the predominant SSO framework in government, education and enterprise environments, while its competitors such as OAuth 2.0 [3] and OpenID Connect (OIDC) [4] are more commonly used on the open Web.
SSO frameworks ease the burden of authentication on SPs, but also impose stricter security requirements for IdPs and the SSO framework itself, as they are a single point of failure — a compromised IdP or an SSO system weakness can lead to compromises of sensitive user data across different SPs. For example, in the Sunburst [5, 6] supply chain attacks, after the initial compromise, the attackers obtained trusted signing certificate used for signing SAML and used it to access all services within the targeted enterprise.
There have been numerous efforts in empirical security analysis of OAuth 2.0 [7, 8, 9, 10, 11, 12], OIDC [13] and SAML [14, 15, 16, 17], resulting in improved protocol implementations and specifications. Whilst the empirical security analysis can identity flaws and improve security, it can not guarantee the absence of weaknesses. On the other hand, formal security analysis can offer valid proofs, sometimes machine verified or even machine generated, that the mathematical model of the protocol holds desired security properties. Of course, the security guarantees provided by a deployed system depend on the faithfulness of the mathematical model to the specification and implementation.
Despite its mission-critical nature and wide adoption, SAML V2.0 has received very little attention from the formal methods community, with the work by Armando et. al [18, 19, 20] being the only attempt to formally analyze systems based on the SAML framework. Their research, however, is exclusively focused on the so called Redirect/POST use case which is only one of several patterns for building Web SSO with SAML (albeit the most common one in use today). Furthermore, Armando et al. only considered one specific deployment configuration used by Google, while SAML standard allows for multiple protocol variants and features. There are no attempts at formal analysis of other use cases, let alone to model the entire specification.
One of the obstacles for formal analysis is that SAML is not itself a security protocol — it is a set of building blocks that can be combined, configured and implemented in many different ways to build the protocol that achieves the desired SSO functionality. Even when we fix the specific SSO use case (e.g., the Redirect/POST use case is this paper), there are at least two additional sources of flexibility — optional features and under-specified mechanisms. While such flexibility is sometimes purely functional, there are cases where both optional features and under specified mechanisms do have an influence on the resulting security properties.
In this paper, we focus on the SP-initiated SSO with POST/Artifact bindings use case and attempt to model the specification, accounting for the flexibility allowed by the standard. We develop a straightforward methodology that, first, identifies sensitive optional features and sensitive under-specifies mechanisms. Such features and mechanisms can be instantiated in several ways with each instantiation resulting in what we call a protocol variant. An example of optional features are digital signatures of authentication requests, which are recommended (but not required) by the SAML standard. Finally, we perform automated analysis of protocol variants yielding security proofs or falsifications of desired properties.
There are multiple benefits to this approach. Most importantly, we can use automated formal analysis tools to examine the exact impact of each feature on the overall security guarantees provided by the protocol, as SAML standards only gives vague low level guidelines. Using the example of optional signatures, the standard [21] says ”The SAML request MAY be signed, which provides both authentication of the requester and message integrity.” — it is unclear what exactly (if anything) can go wrong in systems configured not to require signed authentication requests. Our analysis gives an answer to this question — for the POST/Artifact use case, adding signatures to authentication requests does not have a significant influence on the security of the use case in the symbolic model.
The analysis presented in this paper, while significantly more comprehensive than the related work [18, 19, 20], still fails short of fully modelling the use case since we only consider three degrees of freedom, and model only two options for each. However, this is the first analysis of an important and widely used SAML Web SSO use case (three EU countries, support for this use case as part of their eID services [22]). Furthermore, the result of the formal analysis validate the methodology, demonstrating that protocol variants do indeed have different security features.
In order to perform the analysis we needed to overcome a number of technical challenges related both to SAML itself and modeling with the Tamarin prover. For example, we needed to examine various SAML implementations to identify how exactly are under-specified mechanisms implemented in practice. Also, we carefully model trust establishment steps to account for arbitrary complex system and trust relationships. Finally, we develop a novel method of modelling TLS communication with Tamarin prover since we found existing approaches unsuitable for this purpose.
In summary, the contributions of this paper are as follows:
  • We build the first formal model of the SAML 2.0 POST/Artifact use case. The model is based on the SAML standard rather than a particular deployment, and created using the methodology developed in this paper. The result is a meta-model of the POST/Artifact use case, that we instantiate to eight different protocol variants.
  • While the developed methodology is simple and straightforward, we consider it an independent contribution and an important first step towards comprehensively modeling and analyzing SAML Web SSO profile in its entirety.
  • We give automated formal proofs of the most important security properties of the eight protocol variants including multiple authentication, secrecy and freshness properties. The properties hold in a general setting with an unbounded number of protocol participants and sessions, and a Dolev-Yao adversary that can corrupt arbitrary participants.
  • We identify a weakness in all protocol variants — the artifact-session confusion attack. While the discovered attack does not seem to have an impact on the security of deployed systems, it highlights the necessity of closely following the SAML assertion processing rules.
  • We include the Tamarin code and the scripts needed to reproduce the results as a submission artifact and plan to make the complete model publicly available upon paper publication.
Organization of the paper
Brief description of the SAML standard and the SAML Web SSO Profile is given in Section II. Then, we present the methodology and the formal model of the POST/Artifact use case in Section III. In Section IV, we present the security properties and the result of their verification with the discussion summarising the results in Section V. We discuss related work is Section VI and conclude in Section VII.
II Background
Figure 1: An example protocol execution of SAML V2.0 SSO SP-Initiated SSO with POST/Artifact bindings use case.
II-A SAML Framework
The Security Assertion Markup Language version 2.0 (in the rest of paper abbreviation SAML will denote SAML version 2.0) is a technical framework for exchanging security information between partners [23]. It is an open standard developed by the Security Services Technical Committee under OASIS. The brief technical overview of the SAML standard is given in the official document [23], whereas the standard itself is defined in a series of technical documents [21, 24, 25].
SAML framework defines the following basic building blocks: assertions, protocols, bindings and profiles. Assertions define the syntax, semantics and processing rules for security statements. Protocols define syntax, semantics and high-level processing rules for requesting and delivering assertions. Specification for both assertions and protocols includes various application-level cryptographic mechanisms (e.g., digital signatures for ensuring authenticity of authentication requests, public-key encryption for protecting the confidentiality of assertion data). Bindings map the abstract protocols to specific delivery mechanisms (e.g., SOAP, HTTP Redirect, HTTP POST, artifact). The bindings specification also mandates a transport layer security mechanism — the TLS protocol — in certain situations. Finally, profiles combine SAML assertions, protocols and bindings with an additional set of rules in a system that aims to provide a specific functionality (e.g., a SSO mechanism on the web).
Furthermore, multiple SAML profiles can be configured and combined in different ways to implement specific identity federation SAML use cases [23]. For example, one common use case is SP-Initiated SSO with Redirect/POST Bindings (the “Redirect/POST” use case, in short). Finally, each SAML deployment for a specific SAML use case requires further configuration on all levels (assertions, protocols, bindings and profiles), trust establishment between parties and implementation of system parts that are left unspecified by the SAML framework (e.g., the authentication of end users).
II-B Web Browser SSO Profile
In the SAML Web Browser SSO Profile, an end user uses a web browser client to access a resource at a service provider (SP) while verifying its identity to an identity provider (IdP). On a high level of abstraction, when the user attempts to access a resource, the SP hosting the resource generates and communicates an authentication request to an IdP. The IdP, after verifying the end user identity, communicates a response back to SP containing an assertion describing the identity of the end user as well as other metadata. This profile assumes that SPs and IdPs have established trust and that IdPs have methods of authenticating end users. However, the details of trust establishment and user management/authentication are outside the scope of the SAML standard.
The communication of authentication requests and responses depends on the protocol bindings used. For example, in the Redirect/POST use case, an HTTP Redirect (over TLS) is used to deliver the authentication request from the SP to the IdP via the web browser client, whereas the HTTP POST mechanism is used to deliver back the response, again via the web browser client.
With HTTP artifact binding the requests and/or responses can be communicated by reference using so called artifacts. In that case, another SAML profile — the Artifact Resolution Profile — is additionally used to resolve the artifact into an actual request and/or response.
II-C SP-initiated SSO with POST/Artifact Bindings
In this paper, we model and analyze the SP-initiated SSO with POST/Artifact Bindings use case as defined in  [25, 23]. In this use case, the authentication requests are sent using the HTTP POST binding, whereas the responses from IdPs are sent back using the HTTP artifact binding. Furthermore, the flows are initiated by an end user using a web browser client to request a resource at SP (SAML also supports IdP initiated SSO). Note that, from the symbolic protocol analysis perspective, HTTP Redirect and POST delivery mechanisms are equivalent. Still, we will refer to the use case as “POST/Artifact” in the rest of the paper to be consistent with the SAML conventions.
An example execution flow is given in Fig. 1. End user through the web browser client attempts to access a resource at SP. SP responds with script that initiates a HTTP POST request to the IdP containing the <AuthnRequest> structure — a SAML message consisting of, amongst other data, a unique id, SP identifier and IdP identifier. This <AuthnRequest> can be accompanied with a RelayState message which holds information about end users initial request (this is more discussed in subsection III-A). IdP receives the <AuthnRequest> and tries to authenticate the end user. After successful authentication, IdP builds the SAML <Response> message, but does not deliver it via the web browser client. Instead, it creates an <Artifact> SAML message containing an unique ID associated with the specific <Response> message. This <Artifact> is delivered to the SP via the web browser using either HTTP Redirect or POST mechanism. After receiving the <Artifact>, SP builds the <ArtifactResolve> message and sends it directly using SOAP binding to IdP. After successfully verifying the <ArtifactResolve> request, IdP responds with a <Response> containing an assertion. Finally, the SP verifies the assertion and provides the requested resource to the end user.
II-D Tamarin prover
The Tamarin prover [26, 27, 28] is state-of-the-art tool for symbolic modeling and automated analysis of security protocols. It was successfully used to analyse, break, and improve many real-world security protocols such as TLS 1.3 [29], authentication protocols for 5G networks [30], and WPA2 [31].
In Tamarin, protocols are modeled by multiset-rewriting rules, each representing an atomic action. A rule is of the form $id{:}\mathrel{[l]}\mathrel{\mkern-3.5mu}\mathrel{-}\mathrel{\mkern-3.5mu}\mathrel{[a]}\mathrel{\mkern-3.5mu}\rightarrow\mathrel{\mkern-3.5mu}\mathrel{[r]}$, where $l,a,r$ are sets of factspre-conditions, action facts and post-conditions, respectively. In a typical rule, pre-conditions will correspond to a message an agent receives and its internal state, whereas the post-conditions will correspond to the messages the agent sends and the changes to its internal state.
Security properties are most often specified using guarded first-order logic formulas involving actions facts, that are interpreted over execution traces. Tamarin also supports equivalence properties, but we use trace properties exclusively in this paper. The verification procedure tries to automatically verify or falsify the formula in a constraint system containing user-defined rules, built-in adversary rules and model restrictions. The automated procedure can be further refined by the user by supplying a proof-guiding oracle — a sort of heuristics for optimizing automated proving. The property, if verified, holds in a general setting with an unbounded number of protocol participants and sessions unless, of course, different assumptions are explicitly build into the model or the properties.
III Formal model
Now, we present the formal model of the POST/Artifact use case as defined in  [25, 23] (code available at [32]). To obtain the formal model of the system, we need to model the behaviour of the participants defined by the standards — namely the SPs and IdPs, but also the end users participating in the protocol by using a web browser client. Furthermore, we need to explicitly model the trust establishment steps between protocol participants that are implicit in the SSO use case. Since the use case assumes TLS connections which are not provided by default in the Tamarin prover, we build primitives that model web request-response communication over one-way authenticated TLS. Finally, we need to model the adversary that, besides the usual Dolev-Yao capabilities, can corrupt other protocol participants (SPs, IdPs, and clients) and communicate over TLS with honest participants.
III-A Methodology and Model Variants
As mentioned before, there are several degrees of freedom in the SAML specifications, and here we describe the methodology by which we converted the specifications to the the core formal model of SAML and sell as the eight protocol variants we consider in our analysis.
First, as expected from the complex standard aiming to capture different use cases, there are many optional features. These are reflected as optional fields in SAML assertions, or optional steps in SAML protocols. For example, the IdP may include the optional <Conditions> field in the issued assertion. Note, however, if such a field is present, the assertion consumer must validate the condition before accepting the assertion.
Our approach for optional features was as follows:
  • All required fields and processing rules (i.e., those described with MUST key word in the standard) are included in the core model, with he following exception: If the lack of processing rules gives strong evidence that the field plays a purely functional rather than security role, we leave it out of the model.
  • No optional fields and processing rules (i.e., those described with SHOULD or MAY key words in the standard) are included in the core model. However, if the specification processing rules, or our judgment, indicate that the optional fields could potentially have an impact on the resulting security properties, we include the optional field in the list of sensitive optional features.
Example required fields not included in the model are the <Version> and the <Timestamp> of the <Assertion>. An example sensitive optional feature is the use of <EncryptedAssertion> instead of <Assertion> in authentication responses. This feature is clearly security relevant and we recognize as such. Note, however, we leave the implementation for future work rather than adding it to the core model right now. Table I contains the exhaustive list of all fields of <AuthnRequest> and <Response> not included in our core model with the fields we consider security sensitive.
SAML element full path Required Security sensitive Element description
¡AuthnRequest¿  ¡Subject¿
Specifies the subject of the authentication request or assertion, which identifies the principal that is the subject of the request or assertion.
¡AuthnRequest¿  ¡Version¿
Specifies the version of the SAML protocol.
¡AuthnRequest¿  ¡IssueInstant¿
Specifies time at which the request or assertion was issued.
¡AuthnRequest¿  ¡ForceAuthn¿
Indicates whether the identity provider must authenticate the principal even if the principal has an existing session with the identity provider.
¡AuthnRequest¿  ¡IsPassive¿
Indicates whether the identity provider must not interact with the principal, such as by prompting the principal for credentials, during the authentication process.
¡Response¿  ¡EncryptedAssertion¿
Represents an assertion that has been encrypted to protect its confidentiality.
¡Response¿  ¡Assertion¿   ¡AuthnStatement¿
Contains information about the authentication of a principal, including the time of authentication and the method of authentication used.
¡Response¿  ¡Assertion¿   ¡Subject¿    ¡SubjectConfirmationData¿     ¡OneTimeUse¿
Indicates whether the authentication request or assertion can only be used once, to prevent replay attacks.
¡Response¿  ¡Assertion¿   ¡Subject¿    ¡SubjectConfirmationData¿     ¡NotOnOrAfter¿
Specifies the time after which the authentication request or assertion is no longer valid.
Legend:
 Second column indicates whether the element is required by SAML standard or not.
 Third column indicates whether the element has an impact to the security properties by our judgment.
TABLE I: Elements that are not implemented in our model which are required by SAML standard or/and relevant to the security properties of protocol.
The other source of difficulty when converting the standard to the formal model is the lack of specification of various mechanism. Here, we tried to asses if the under-specified mechanism plays a purely functional role or if it could potentially have an impact on the security properties of the overall system. In the latter case, we include the mechanism in the list of sensitive under-specified mechanisms. For example, the SAML standard specifies that “dereferencing of the artifact using the Artifact Resolution profile MUST be mutually authenticated, integrity protected, and confidential“ [25], but does not explicitly specify the mechanisms to achieve these properties.
When deciding how exactly to model an under-specified mechanism, we considered a straightforward canonical implementation and, additionally examined various SAML implementations to identify how are those under-specified mechanisms implemented in practice.
After identifying the sensitive optional features and sensitive under-specified mechanisms, we decided to pick three of these and model each in two different ways yielding eight different protocol variants. Since our goal was to investigate if and how small changes impact the security of the complete system, we picked the features and implementations that seemed most likely to have a significant impact, but are also illustrative of kinds choices and potential pitfalls in building a standards-compliant SAML deployment.
Technically, protocol variants are implemented by building a meta model with simple preprocessor directives enabling us to automatically generate eight models for all combinations of selected features.
Authentication request signing
As mentioned before, signatures in <AuthnRequest> messages are optional in the SAML standard. Informally, signatures guarantee the integrity and authenticity of the <AuthnRequest> messages and may be included in deployments where the authentication of the request issuer is required. Hence, we created the “signed” and “unsigned” model variants with signed and unsigned <AuthnRequest> messages in order to precisely determine the impact of those signatures.
RelayState mechanism
Some SAML bindings allow the use of the RelayState mechanism to preserve state information. In a POST/Artifact use case, the SP may include the arbitrary RelayState data alongside the <AuthnRequest> — the IdP must return the same data along with the protocol response. Typically, RelayState data enables the SP to “look up” or reconstruct the original request and take the appropriate action upon receiving the assertion. We consider the RelayState mechanism a sensitive under-specified mechanism since it could potentially impact the resource authenticity property — informally, the resource the client is delivered has to be the same resource it originally requested.
We model two different behaviors corresponding to the two different options for the RelayState mechanism in the popular open source SAML framework Shibboleth [33]:
  • In the ”norelaystate” model variant, no state is stored in the SP and the URL requested by the client is sent as RelayState data. When the SP receives the response (the <Artifact> message and the RelayState message containing the URL), it assumes that the URL is the original URL requested by the client. This behavior is consistent with the default behavior of Shibboleth Service Provider v3.
  • In the ”relaystate” model variants, the SP generates a new nonce, stores it in the session data along with the original URL, and inserts the nonce into the RelayState data. This nonce is used to join the original SP session when the response is received. This behavior is equivalent to the ”ss:mem” value for the Shibboleth Service Provider v3 ”relayState” option (which is also the recommended option).
SAML identifiers
SAML specification mandates the use identifiers for many SAML messages. The specification requires that “any party that assigns an identifier MUST ensure that there is negligible probability any other party will accidentally assign the same identifier to a different data object” [21]. Notice that this does not imply that identifiers must be generated via a cryptographic pseudorandom number generator (PRNG) — only that the probability of collisions between identifiers generated by honest protocol parties is negligibly low. Clearly, this is a sensitive under-specified mechanisms since SAML identifiers are used to match responses to the original requests.
We examined several popular SAML frameworks (Shibboleth, opensaml, LemonLDAP, OneLogin) and identified two cases (LemonLDAP and OneLogin PHP SDK) where a general purpose PRNG is used to generate SAML identifiers and, hence, the identifiers can be predicted by the adversary given enough past information.
We model two different behaviours corresponding to cryptographically strong and weak random identifiers:
  • In the “strongid” model variants, the SAML identifiers are fresh nonces and, hence, unpredictable by the adversary.
  • In the “weakid” model variants, the SAML identifiers are fresh nonces that are leaked to the adversary immediately upon generation (and before their use in the protocol). This corresponds the pedantic (but pessimistic) interpretation of the SAML standard — no two honest parties will ever generate the same identifier, but, nevertheless, the adversary knows all the identifiers beforehand.
III-B Communication channels
SAML SSO profiles mandate the use of TLS channels while sending authentication request and responses. Somewhat surprisingly (given the large number of case studies using the Tamarin prover), we found the previous approaches for modelling TLS communication in Tamarin lacking for our purpose and developed a simple novel approach for modelling the “request-response over one-way authenticated TLS” pattern in Tamarin.
Most common approach to modelling protected communication is by using secure channels as in [34]. However, we found these to be too imprecise for our purpose. Secure channels provide both authenticity and confidentiality, but they authenticate both parties (unlike one-way authenticated TLS used by SAML SSO). Moreover, secure channels do not provide the link between requests and corresponding responses. A more suitable formalization is found in [35] — the TLS sessions are created using a dedicated rule, messages are transferred through facts, and channel IDs are used to bind messages to specific TLS channels. While this model of TLS communication is precise, we found it to be computationally unfeasible for our purpose. The multiplicity of honest agent’s interactions resulted in huge Tamarin precomputation times.
Our model of TLS communication — given in Listing 1 — is similar to one in [35], except that we use ordinary network channels for communication and wrap TLS messages inside private function names. Using ordinary network channels significantly reduced precomputation time while wrapping TLS messages inside private functions makes it impossible for adversary to read the messages. Furthermore, we do not require channel IDs as in [35] since TLS sessions in SAML SSO are short-lived (only one request and response).
1functions: https_request/3 [private],
2           https_response/2 [private]
3rule Client_Starts_TLS:
4[
5  Fr(~tls_session_id),
6  !Is_Agent(\$Client),
7  !Is_Agent(\$Server)
8]
9--[
10  Client_Starts_TLS(\$Client, \$Server, ~tls_session_id)
11]->
12[
13  New_Client_TLS_Session(
14      \$Client,\$Server, ~tls_session_id),
15  New_Server_TLS_Session(\$Server, ~tls_session_id)
16]
17
18rule Sample_TLS_Request:
19[
20  New_Client_TLS_Session(
21      \$Client, \$Server, ~tls_session_id)
22]
23--[]->
24[
25  Out(https_request(\$Server, ~tls_session_id, request))
26]
27
28rule Sample_TLS_Response:
29[
30  New_Server_TLS_Session(\$Server, ~tls_session_id),
31  In(https_request(\$Server, ~tls_session_id, request))
32]
33--[]->
34[
35  Out(https_response(~tls_session_id, response))
36]
Listing 1: TLS Channel rules
TLS session is established between two agents (the notion of agents is precisely described in the following section III-C). For every TLS request towards agent acting as server S, a new session is created and labeled with unique tls_session_id. Since we model one-way authenticated TLS, any agent can act as client and initiate such session. However, only agent S can access the data in the request, and create a response bound to the same tls_session_id Such response can be accessed only by the agent that initiated the TLS session. Note that the agents name S doubles as its TLS identity, in practice this can be, for example, its common name or a pinned certificate.
III-C System setup
[…]
[t](botskip=0pt, midskip=0pt)[scale=0.75]fig/hartl2.pdf Prerequisites Rules state machine. Initialized agent can become one of three honest parties or become compromised. Clients (and the adversary) can register resources at SPs (honest or compromised). Client (and the adversary) can register at IdPs (honest or compromised). Finally, any two SAML parties (SP and IdP) can establish trust. In our model, the system consists of a number of agents. There are three kinds of honest agents – the SPs and IdPs whose behaviour is defined by the SAML specification (the SAML parties) and the end user using a web browser client (the client). Additionally, adversary can create and control any number of compromised agents. In fact, since honest agents only communicate over TLS, the adversary must create and use compromised agents to interact with the system as described in Section III-B.
Each agent in our model is assigned a unique name — a public value in Tamarin. For SAML parties, the agent name plays multiple roles — we use as its domain name in the TLS protocol, as the URL at which the agent receives SAML requests, and as its SAML entity id (a globally unique name for a SAML party). For clients, the agent name is purely a feature of the formal model, and each client is assigned a unique nonce clientID during system setup.
System setup rules model the agent initialization, key generation, role assignment, resource registration, and trust establishment. The rules are general and for allow arbitrary complex system and trust relationships. A state machine illustrating the system setup rules is given in Fig. III-C. We assume unbounded number of agents (SPs, IdPs, Clients) as well as an unbounded of parallel sessions for each agent.
III-D SAML parties – the SP and the IdP
Using the described methodology, building the formal model of the SP and the IdP to match the POST/Artifact use case is a straightforward but tedious process of translating SAML messages and processing rules into Tamarin rules.
As an example, in Listing 2 we provide the rule modelling the behavior of SP when receiving a resource request from an honest agent or adversary.
Client creates new TLS connection with unique tls_session_id and sends requests for resource located at url at $SP$. On the other side, SP awaits such request if other facts are already created:
  • !SP_Initialized(\$SP) — SP is initialized.
  • Is_Resource(\$url, \$SP) — resource was registered during system setup (possibly by the adversary).
  • !Is_Registered(\$SP, \$IdP) — IdP and SP have established trust during system setup.
  • Fr(~rs) — fresh variable denoting relay state is created in this rule step.
  • New_Id(~id_AuthnRequestuest) — precondition modelling creation of a fresh variable for to use as the message id of Authentication Request.
  • New_Server_TLS_Session(
    \$SP, ~tls_session_id_C_SP)
    - newly created TLS channel between C and SP.
Once SP receives the ResourceRequest and all the above facts are available, the SP is creates an authentication request <AuthnRequest>, and sends it along with RelayState back to client.
As mentioned in section III-A, <AuthnRequest>, RelayState, and the rule creating the the term containing the message id — New_Id(~id_AuthnRequestuest) — vary according to the protocol variant. For example in the “strongid” variants the term is created by the rule that keeps the ~id_AuthnRequestuest secret, while and in “weakid” variants ~id_AuthnRequestuest is sent to network and, hence, made known to adversary.
1rule SP_Resource_Request:
2  let
3    // Received
4    ResourceRequest = <’ResourceRequest’, \$url>
5    Req = https_request(
6        \$SP, ~tls_session_id_C_SP, ResourceRequest)
7    // Building
8    AuthnRequestuest = <’AuthnRequest’,
9        ~id_AuthnRequestuest, \$SP, \$IdP>
10    OptionalSignature =
11        <#AUTHN_SIGNATURE AuthnRequestuest | ~ltkSP>
12    RelayState = <#RELAY_STATE_MESSAGE ~rs | \$url>
13    // Sending
14    Resp = https_response(~tls_session_id_C_SP,
15        <\$IdP, RelayState, AuthnRequestuest,
16            OptionalSignature>)
17  in
18  [
19    !SP_Initialized(\$SP),
20    !Is_Resource(\$url, \$SP),
21    !Is_Registered(\$SP, \$IdP),
22    !Ltk(\$SP, ~ltkSP),
23    New_Server_TLS_Session(\$SP, ~tls_session_id_C_SP),
24    In(Req),
25    New_Id(~id_AuthnRequestuest),
26    Fr(~rs)
27  ]
28  --[
29    Start_SP(\$SP),
30    SP_Resource_Request(\$SP, \$IdP, ~id_AuthnRequestuest,
31        \$url),
32    Https_Response(Resp)
33  ]->
34  [
35    SP_State_1(\$SP, \$IdP, AuthnRequestuest,
36        <#SP_SESSION_STATE ~rs | \$url>),
37    Out(Resp)
38  ]
Listing 2: Example Tamarin prover rule - SP recieves resource request
III-E Client modeling
In our work, the focus has been on thoroughly modelling the system parts defined by the SAML technical standards — namely the SPs and IdPs. Since behaviour of the user and web client is not defined by SAML technical standards, we build the canonical user and the web client. We combine the end user and the web browser client into a single entity that we refer to as the client in our formal model and in this paper.
The client is an idealized version of the the end user and it’s web browser client that does not take into account all the intricacies of the browser behaviour. For example, our client will not follow arbitrary redirects, will not store cookies etc. Also, in a more comprehensive model, one would need to take into account end user behaviour and possible human errors. For example, will the end user diligently verify the description of the service provider before authenticated and giving consent at the identity provider.
In our formal model, the client can start the protocol flow with any service provider requesting any resource. Hence, client can initiate sessions even with service providers it is not registered with and even with resources belonging to other clients. This models possible mistakes or social engineering attacks on the end user.
During authentication to the IdP we assume that the end user will authenticate and consent only when the SP matches the service provider the client initiated the session with. Hence, we assume that the user is diligent enough to recognize the intended SP.
Moreover, the client will remember the SP it initialed the protocol with and only again connect to that specific SP when it received the artifact from the SP. In the real world, this is implemented using a combination of relay state and browser cookies.
Finally, client authentication to the IdP is also not part of the SAML standard, so we wrote our rules for client authentication. Before authentication occurs, client needs to be registered at IdP for witch we also wrote custom rules described in section III-C. Client has shared his $clientSecret$ (representing clients credentials) with IdP. While authenticating, after providing $clientSecret$ to the IdP over secure channel and after IdP successfully verifies the clients credentials, client is considered authenticated and IdP can now issue authentication assertion.
III-F The Adversary
As mentioned before, HTTPS requests and responses which are modeled by private functions and, hence, the adversary cannot construct them directly. Therefore, we introduce special rules that enable the adversary to send and receive HTTPS requests and responses, acting either as a server and as a client.
To participate in protocol flow, the adversary first needs to compromise some of honest agents and then he can communicate in their name with other honest agents. Two compromised agents are not able to communicate to each other. By implementing the adversary this way we enabled Tamarin built-in engine for the adversary which can create fresh facts, combine facts etc. Initialization rules that enable the adversary to pose for interaction with the protocol participants are marked with red arrows on Fig. III-C.
III-G Limitations
We make a number of simplifying assumptions to keep the complexity of the model manageable. We encountered some parts that are difficult to model in formal fashion or make the verification computationally unfeasible. Moreover, we needed to add some additional parameters for executability in formal model and for expressing security properties.
Client Fidelity
Clients are a more idealized than SPs and IdPs as described in III-E.
Client Type Checking
Rules for client communication with SP and IdP are basically describing the behaviour of a typical web browser. Aside from the initial request, other two client rules are just automatic web browser responses to HTTP scripts that are embedded as SPs and IdPs requests and responses. This could have been implemented in formal model by just passing the message from the input to the output, but because of formal model limitations, we implemented type checking of this automatic HTTP script requests and responses. Here we carefully designed rules so the client is not comparing values of message variables, just checking variable type (fresh, public etc.). Using type checking we removed significant number of meaningless message exchange through client automatic rules and therefore reduced Tamarin reasoning time. Of course, such type checking is not used in real code implementations and our model is, therefore, less faithful in that regard. Nevertheless, we argue that type checks are not a significant loss of generality, since the type checking is definitely happening at SP and IdP side. Moreover, for cases where the adversary would like to send something that does not correspond to message type, it can still do that through the adversary rules.
SAML SSO Profile known messages fields
SAML SSO Profile defines number of properties and message definitions, message ids, name-spaces, additional assertion parameters etc. These fields are well-specified in the protocol documentation and are not subject to modification or tampering by adversary. These properties are important in concrete implementations, but on formal level of abstraction they are irrelevant.
SAML SSO Profile additional execution paths
Modeling too many execution paths can make the analysis intractable and the results difficult to interpret. Therefore we focused on the critical aspects and security-relevant behaviors of the POST/Artifact use case and did not model all possible execution paths in SAML SSO Profile that we modeled. Example execution paths that are not modeled are if client sends wrong request then SP returns error, if SP sends wrong format authentication request then IdP returns error, if client sends multiple resource request etc.
Expressing security properties
For expressing security property of resource secrecy we needed to add additional parameters out of scope of SAML specification. While delivering resource to Client in final step, SP generates fresh unique variable which we called temporary secret $ts$. Without this session unique value we could not express property regarding the secrecy of resource delivered to client in specific session. We did the same with assertion issued by IdP to SP where we inserted assertion secret variable $sec$.
IV Security properties and analysis results
In this section, we describe the security properties we consider and discuss the results of the formal analysis.
No. Property Protocol variant
srs srw sns snw urs urw uns unw
1 sec_Client_Resource_Authenticity
2 sec_Client_Resource_Authenticity_Strong
3 sec_Client_Registered_Resource_Secrecy
4 sec_Client_Registered_Resource_Secrecy_Strong
5 sec_Client_Resource_Secrecy
6 sec_Client_Resource_Secrecy_Strong
7 sec_Client_Resource_Freshness
8 sec_Client_SP_Non_Injective_Agreement
9 sec_Client_SP_Non_Injective_Agreement_Strong
10 sec_Client_SP_Injective_Agreement
11 sec_Client_SP_Injective_Agreement_Strong
12 sec_SP_Client_Resource_Secrecy
13 sec_SP_Client_Resource_Secrecy_Strong
14 sec_SP_Client_Non_Injective_Agreement
15 sec_SP_Client_Non_Injective_Agreement_Strong
16 sec_SP_IdP_Weak_Agreement
17 sec_SP_IdP_Weak_Agreement_Strong
18 sec_SP_IdP_Non_Injective_Agreement
19 sec_SP_IdP_Non_Injective_Agreement_Strong
20 sec_SP_IdP_Authentication_Freshness
21 sec_SP_IdP_Authentication_Freshness_Strong
22 sec_SP_IdP_Assertion_Secrecy
23 sec_SP_IdP_Assertion_Secrecy_Strong
24 sec_IdP_SP_Non_Injective_Pre_Agreement
25 sec_IdP_SP_Weak_Agreement
26 sec_IdP_SP_Non_Injective_Agreement
27 sec_IdP_SP_Injective_Agreement
28 sec_IdP_SP_Assertion_Secrecy
Legend:
 first letter of the protocol variant – authorization requests are signed (“s”) or unsigned (“u”)
 second letter of the protocol variant – relay state mechanism is used (“r”) or not used (“n”)
 third letter of the protocol variant – random IDs are cryptographically strong (“s”) or weak (“w”)
– property verified, – property falsified
TABLE II: Analysis results for each property and protocol variant.
The security properties can be roughly grouped into authentication properties, secrecy properties and freshness properties. We consider all variants of authentication properties described in the seminal paper [36] — namely weak, non-injective and injective agreement. Secrecy properties include secrecy of assertions and resources. Finally, we consider freshness properties of particular protocol parameters. Properties can be, furthermore, grouped by the participant from whose point of view we consider the protocol execution (client, SP and IdP), and by the exact assumptions on entities controlled by the adversary.
It is important to note that the desired security properties are not explicit or precise in the official SAML documentation. The said documentation includes the “Security considerations” non-normative document [37] containing high-level descriptions of the desired privacy and security properties as well as the threat model. Additionally, it gives multiple practical guidelines for implementing systems based on the SAML framework in the secure manner. In comparison with SAML documentation, we consider a more comprehensive set of security properties, and provide their precise formalization.
In total, we have considered 28 security properties and have either formally verified or falsified every property for each of the 8 protocol variants. The complete overview of the verification results is given in Table II, and the rest of this section gives an overview of the most relevant properties.
IV-A Client PoV
First, we investigate the security properties from the client’s point of view. That is, we assume that a client that has successfully completed a protocol session apparently with the SP $S$ and the IdP $I$, and investigate what guarantees hold for the client under different assumptions on honesty of other protocol principals. The honesty assumptions we consider are as follows:
  • Both the service provider $S$ and the identity provider $I$ used in this particular session are honest.
  • The service provider $S$ used in this particular session is honest, and additionally the service provider $S$ has only established trust with honest identity providers.
Hence, whenever it made sense, we considered two versions of security lemmas: the strong and the weak version corresponding to the first and second assumption above. By considering the two versions, we are aiming to measure the damage to the protocols in scenarios where a service provider establishes trust with an identity provider which happens to be compromised.
Client resource authenticity
If a client $C$ completes a protocol session with the SP $S$ and the IdP $I$ with session parameters client ID $cid$, obtaining resource $res$ at URL $url$ then $C$ has registered that exact resource $res$ located at $url$ at the SP $S$.
As mentioned above, we also consider two versions of property — the weak and the strong version. In the stronger version, instead of assuming that $S$ does not trust any compromised identity provider $I^{\prime}$, we only assume that the specific identity provider $I$ used in this particular session is not compromised. Let us elaborate a bit more: assume a deployment where the service provider $S$ has established trust with three different identity providers $I_{1}$ , $I_{2}$ and $I_{3}$ such that $I_{3}$ is compromised while the other two identity providers are honest. If the strong version of client resource authenticity holds, that means that the resource provided in a particular session is authentic as long as an honest IdP (e.g., either $I_{1}$ or $I_{2}$ was used in that particular session). If only the weaker variant of the property holds (and stronger does not), that means that the client resource authenticity cannot be guaranteed in this deployment – even if an honest IdP is used in a particular session, the adversary could still violate the resource authenticity property by launching an attack (that somehow takes advantage of the fact that IdP $I_{3}$ is dishonest).
The formal analysis shows that this security property holds for both weak and strong version for all tested protocol variants. Hence, in the SP-initiated SSO with POST/Artifact use case, the client is guaranteed that the resource he is provided with is the resource that it registered as long as the SP and IdP used in the session are both honest.
Client registered resource secrecy
If a client $C$ registers a resource $res$ at the SP $S$ then that resource remains forever unknown to the adversary.
In the strong variant of property, we expect the secrecy to hold even if there exists some compromised IdP $I^{\prime}$ trusted by $SP$. This variant is included for sanity checking purposes as it should clearly fail — a compromised IdP $I^{\prime}$ trusted by $S$ can initiate a session, craft authentication responses for any client of its choice, and obtain their resources.
The formal analysis shows that the weaker version of the property holds for all protocols considered, whereas the strong version fails for all protocols variants as expected.
Note that, this property considers the secrecy of the long-lived resources across sessions. The following property considers the resource secrecy on the session level.
Client resource secrecy
If a client $C$ completes a protocol session with the SP $S$ and IdP $I$ with parameters client ID $cid$, and resource $res$ located at URL $url$ accompanied with temporary secret $ts$, then the temporary secret $ts$ remains forever unknown to the adversary.
Temporary secret $ts$ is of type fresh and is created and delivered in last step when SP delivers resource to client. This property means that if client manages to complete the protocol and to acquire resource in that concrete protocol run, it is not possible that the adversary obtained the resource compromising that particular session.
The formal analysis shows that this security property holds for both weak and strong version for all protocol variants.
Client resource freshness
If a client $C$ with ID $cid$ completes the protocol session with the SP $S$ obtaining resource $res$ accompanied with temporary secret $ts$, then the temporary secret $ts$ was created fresh after $C$ has requested the resource. This property holds unless $S$ is compromised.
Informally this property states that the client cannot be tricked into accepting an old resource. This property does not depend on the identity provider and, hence, we do not make any assumptions on the honesty of IdPs present in the system.
The formal analysis shows that this security property holds for all protocol variants.
Client – SP non-injective agreement
If a client $C$ with ID $cid$ completes the protocol session with the SP $S$ and IdP $I$ obtaining resource $res$ at $url$ with temporary secret $ts$, then $S$ has completed a protocol with $cid$ and the session parameters (the resource $res$, URL $url$ and temporary secret $ts$) all match.
This is a standard two party authentication property — if a client has $C$ completed a session with $S$ then there is a session of $S$ where the session parameters agree with those in the client session.
The formal analysis shows that this security property holds for both weak and strong version for all protocol variants.
The following property requires that, additionally, there are no multiple sessions with the same session parameters.
Client – SP injective agreement
If a client $C$ with id $cid$ completes the protocol session with the SP $S$ and IdP $I$ obtaining resource $res$ at URL $url$ with temporary secret $ts$ then $S$ has completed a protocol with client represented by $cid$ and the session parameters match and, additionally, there must not exist a client session of with same parameters $res$ and $cid$.
The formal analysis shows that this security property holds for both weak and strong version for all protocol variants.
IV-B Service Provider PoV
From the SP’s PoV, we considered similar security properties as for client’s PoV. In addition to authentication properties between the client and SP, we consider authentication properties between SP and IdP. Furthermore, we also consider resource and assertion secrecy. As before, wherever applicable, we consider the weak and the strong version of the property depending on the exact assumption on the honest IdPs.
The analysis shows that the resource secrecy (both weak and strong) and the non-injective agreement (again both weak and strong) between the SP and the client hold for all protocol variants.
It is important to note that these properties between the client and the SP are not redundant or implied by the analogous properties from the PoV of the client. For example, SP – client resource secrecy property says that the temporary secret $ts$ served by the SP $S$ in some session remains unknown to the adversary as long as the client that registered the corresponding resource and the IdP $I^{\prime}$ used in that particular session are not compromised. This property guarantees that the honest client’s resources remain secret even in the scenarios when the client mistakenly establishes or completes sessions with compromised service providers — and these scenarios are not covered by the analogous properties from the PoV of the client. We do not include the SP – client injective agreement in the list of properties since it is trivially implied by the non-injective agreement — the session parameters (e.g., the ID $ida$ and the temporary secret $ts$ are generated fresh by $S$ making the SP’s session trivially unique).
Taken together with verified client – SP injective agreement positive results, we prove that SP-initiated SSO with POST/Artifact binding offers strong bidirectional agreement between the client and the SP on the session level as long as the IdP used in the session is honest.
Next, we consider authentication properties between the SP and the IdP (namely the weak agreement and the non-injective agreement), and show that the weak and strong versions of both properties hold for all protocol variants. The agreement between the SP and the IdP is probably the most important property of the protocol as it captures the purpose of the SSO — if a SP accepts the assertion from an IdP, then IdP must have issued the assertion after authenticating the client. In all properties between SP and IdP we make no assumptions on the honesty of the clients. Hence, the properties hold when all clients are under the control of the adversary.
In the following property descriptions, $ida$ is the unique identifier of an <AuthnRequest> message while $mh$ is an unique message handle of an <Artifact> and $idr$ is unique identifier of an <ArtifactResolve> message.
SP – IdP non-injective agreement
If a SP $S$ completes a protocol with client with ID $cid$ and the IdP $I$, then $I$ had completed a protocol and authenticated the client with ID $cid$. Furthermore, and all other parameters ($mh$, $ida$ and $idr$) match.
Again, injective agreement property is trivially implied by the non-injective agreement is, hence, not included in the analysis. We do, however, consider one additional agreement property that, informally, requires that the IdP has authenticated the client recently.
SP – IdP authentication freshness
In addition to the non-injective agreement above, this property requires that the authentication of the client with ID $cid$ is fresh in sense that it happened after the $S$ has sent the authentication request.
The analysis verifies this properties for 6 protocol variants and falsifies the property for 2 protocol variants. The property does not hold in variants where the authentication requests are not signed and the IDs are cryptographically weak. In the attack that violates this property the adversary guesses the ID used for authentication request and uses it to construct an authentication request that appears to be originating from $S$. The result is that the client authentication is triggered before the genuine authentication request is issued by the SP $S$. We do not consider the violation of this property to be a serious drawback of the protocol in practice as discussed in Section V.
Finally, we consider the secrecy of assertions accepted by the SP. Authentication assertion can possibly hold some sensitive user data, and therefore its secrecy is also important in some deployments.
SP – IdP assertion secrecy
If a SP $S$ accepts an assertion containing an assertion secret $sec$, then that assertion remains unknown to the adversary.
Formal analysis shows that both the strong and the weak version of the property holds in all protocol variants.
IV-C Identity provider POV
From the IdP’s PoV, we consider several authentication properties between IdP and SP as well as the secrecy of assertions issued by the IdP. In all properties we assume that the SP $S$ is honest and, again, make no assumptions on the honesty of clients in the protocol.
IdP – SP non-injective pre-agreement
If an IdP $I$ has completed the protocol and issued an assertion to the SP $S$ in response to an <AuthnRequest> with ID $ida$, then $S$ has issued the <AuthnRequest> with ID $ida$ and has chosen $I$ as the IdP for the session.
Formal analysis verifies this properties for the four protocol variants where the authentication request are required to be signed and falsifies the property for the four protocol variants where the authentication requests are unsigned.
This is not too surprising — if an IdP accepts unsigned authentication requests, then the adversary can issue them arbitrarily. However, the exact failure mode is more subtle since it requires that $I$ completes the protocol, which involves resolving the artifact, which, in turn, requires mutual authentication between the SP $S$ and the IdP $I$.
In the attack falsifying the property, adversary acting as a client obtains a genuine <AuthnRequest> from the SP $S$, but replaces the identifier $ida$ with a different value before forwarding the request to the $I$. The $I$ will, now, complete the protocol — it will resolve the artifact even though there is a mismatch between the $ida$ parameters in the sessions of $I$ and $S$.
Note that the SP $S$ will not complete the protocol session — in the last step it will reject the assertion since the “InResponseTo” field in the assertion does not match the original <AuthnRequest> ID $ida$. This behavior agrees with the assertion processing rules in the SAML specifications. Again, we do not consider this failure to have a significant impact in real world deployments.
IdP – SP weak agreement
If an IdP $I$ has completed the protocol and issued an assertion to the SP $S$, then $S$ has issued the corresponding <AuthnRequest> choosing some $I^{\prime}$ as the identity provider, and also requested the <Artifact> that $I$ resolved.
It is straightforward to conclude that weak agreement should hold due to mutual authentication during the artifact resolution phase. Formal analysis indeed shows that this property holds in all protocol variants.
IdP – SP non-injective agreement
If an IdP $I$ has completed the protocol and issued an assertion to $S$, then $S$ has issued the corresponding <AuthnRequest> choosing $I$ as the identity provider and also requested the <Artifact> that $I$ resolved. Hence $I$ and $S$ agree on the parameters of the <AuthnRequest> and the <Artifact>.
Formal analysis falsifies this property for every protocol variant. This is somewhat surprising at the first glance, but deeper analysis reveals that the property should indeed fail. The fundamental reason is similar to non-injective pre-agreement above — from the IdPs point of view there is no way of verifying that the artifact the SP is attempting to resolve corresponds to the original authorization request.
In the protocol execution falsifying this property, adversary takes the role of a compromised client and perofrms what we call an artifact-session confusion attack as follows:
  1. 1.
    Adversary starts a first session with the SP $S$ requesting the resource $res_{1}$ and obtaining a genuine <AuthnRequest> with ID $ida_{1}$.
  2. 2.
    Adversary starts a second session with the SP $S$ requesting the resource $res_{2}$ and obtaining a genuine <AuthnRequest> with ID $ida_{2}$.
  3. 3.
    Adversary continues the second session, authenticates with IdP $I$ and obtains a genuine <Artifact> with handle $mh_{2}$.
  4. 4.
    Adversary continues the first session and sends the <Artifact> with handle $mh_{2}$ to $S$.
  5. 5.
    SP $S$ resolves the received <Artifact>, causing the $I$ to complete the first session resolving <Artifact> with handle $mh_{2}$ for <AuthnRequest> with ID $ida_{1}$.
The end result of the execution is that IdP has completed a protocol session but there is no session of $S$ that agrees on the parameters of <AuthnRequest> and <Artifact>.
This attack is consistent with SAML specification for the following reasons: First, there is no way for $S$ to check that the <Artifact> received in step 4 corresponds to a different session – the artifact is opaque by design and the only contains the information about its originator $I$. Second, it is neither required or possible for $I$ to perform additional checks during the artifact resolution protocol in step 5. We were also able to confirm this attack is possible in practice by launching it against our simple deployment consiting of Shibboleth [33] Service Provider v3 and Identity Provider v4.
Again, the damage to the protocol is limited due to the fact that the a SAML compliant SP $S$ will not complete the protocol and issue the resource since abort the protocol after detecting the mismatch between the received assertion and the original <AuthnRequest>. Hence, we believe that this attack has no direct impact on the security of deployments, but it highlights the necessity of following the assertion processing rules specified in the SAML standard.
Finally, we consider the secrecy of assertions from the PoV of the IdP.
IdP – SP assertion secrecy
The <Assertion> that IdP $I$ issued to the service provider $S$ remains unknown to the adversary.
Formal analysis verifies this property for every protocol variant. Taken together with the analogous property from the PoV of the SP, we prove that SP-initiated SSO with POST/Artifact binding using case guarantees assertion secrecy even if the assertions are not encrypted.
IV-D Challenges of automated verification
The Tamarin meta model (with comments removed) contains 812 lines of code, about half of which correspond to protocol rules and half to lemmas and security properties. Each of the 8 concrete models for protocol variants is of roughly the same size as the meta model. Even though the models are of moderate size, the analysis proved challenging with many proofs never terminating, even in the simplified setting with a bounded number of sessions. Hence, several insights were needed to achieve termination and to finally verify/falsify all properties in a general setting with an unbounded number of sessions.
One source of complexity were the TLS channel rules initially implemented through facts — similar how TLS channels have been modelled in Tamarin before [35, 38] (which is, in turn, similar to confidential and authentic channels in the Tamarin documentation). This approach, however, resulted in exploding complexity during the precomputation phase due to the enumeration of all possible interactions of honest protocol participants and the adversary via the TLS channels. We addressed this complexity by, first, modelling TLS channels using a combination of private function symbols and adversary rules as described in Section III-B. This, along with a source lemma to resolve partial deconstructions, significantly reduced the complexity of precomputation (while, of course, moving some of the complexity to the main analysis phase). Secondly, we proved a number of TLS-related auxiliary lemmas further reducing the complexity of the analysis.
For most properties, further optimizations were needed. One other source of complexity is the large number of fresh values used in the protocol (in particular, most SAML messages require IDs which are modelled as fresh values). In each of our protocol models, there were roughly 50 refined sources for the $KU(\sim{}t)$ fact which translates to 50 new branches in the search tree each time a goal of such type is processed. We finally tamed the complexity by writing proof-guiding oracles for each problematic lemma, or groups of related lemmas.
Also, we significantly improved the analysis time of the security properties considered from the SP and the IdP PoV by removing all client rules from the model. Since there is no assumptions on the honesty of clients in those properties and, furthermore, all client actions can be performed by the adversary (as evidenced by the executable lemmas), there is no loss of generality here.
Finally, when running sanity checking tests (the executable lemmas) we restricted the model to a bounded number of sessions (only one client, one SP and one IdP) and achieved much better analysis times.
No. Lemmas Protocol variant
srs srw sns snw urs urw uns unw
1 Executable lemmas 121s 121s 120s 104s 126s 116s 101s 98s
2 Helper lemmas 104s 104s 98s 97s 103s 103s 96s 96s
3 Client PoV properties 5813s 5810s 3042s 3035s 6382s 6396s 2976s 2981s
4 SP and IdP PoV properties 425s 431s 403s 403s 1569s 242s 253s 230s
TABLE III: Total analysis times (in seconds) for all lemma groups and protocol variants.
Final analysis was performed using Tamarin version 1.6.1 on an 8-core/16-thread desktop machine with Intel i9-9900K CPU and 32G of RAM. The total analysis time for all properties and protocol variants is close to 10 hours (160 hours of CPU time); the breakdown per protocol variant and the type of lemma is given in Table III.
V Summary and discussion of main results
Based on the results of the analysis we make the following claims on the POST/Artifact use case:
V-A The POST/Artifact use case satisfies desired security goals for a Web SSO system.
Formal analysis verifies the most important security properties for all the considered protocol variants. Some security properties were falsified for some protocol variants, however, these failures are either expected or do not pose a significant threat in the real-world deployments of the use case.
The violation of the SP-IdP Authentication Freshness property in some protocol variants, means that the IdP can authenticate clients before the SP requests their authentication. However, SAML assertions contain additional attributes that can, and should, be used when the time of authentication is of importance. For example, the <AuthnInstant> attribute specifying the time at which the authentication took place or the <NotBefore>/<NotOnOrAfter> attributes specifying the validity time period of the assertion.
The artifact-session confusion attack breaking the IdP-SP non-injective agreement property is interesting, but the extent of the attack is that an IdP can be made to resolve authentication responses that SP will never consume. Hence, we believe that there is no practical impact of this attack, but repeat that care must be taken by the SP to follow the assertion processing rules completely.
Additionally, the use case provides a certain level of robustness against a compromised IdP. Of course, a compromised IdP may still impersonate and access resources of any user at any trusting SP (as we demonstrate by falsifying the strong version of the Client registered resource secrecy property). However, a compromised IdP (even if trusted by SPs) cannot degrade security of protocol sessions where an honest IdP is used.
V-B Details left unspecified by the SAML standard can have an impact on the exact security properties provided by the use case.
The protocols with cryptographically strong IDs have slightly different security properties then those with weak IDs. Admittedly, the difference is subtle, only present when combined with unsigned authentication requests, and unlikely to have any practical consequences. However, we believe the results validate the general concern that the seemingly purely functional choices left to developers by the SAML standard can end up having an impact to the security of the implemented system. However, the formal analysis did not discover any difference between the two different relay state mechanisms considered.
V-C There is no clear mapping between SAML optional protection mechanisms and provided security guarantees.
Adding signatures to authentication requests does not have a significant influence on the security of the use case. For the most important security properties there is no difference between the variants where signatures are required and when they are not. Besides the above mentioned authentication freshness issue, the only difference is the IdP-SP non-injective pre-agreement property where the adversary can make the IdP issue unsolicited authentication responses (that the SP will ultimatelly reject). This claim should be taken only in the limited context of symbolic protocol analysis — in a real-world deployment an IdP that requires signed authentication requests may be much less vulnerable to various low-level attacks (compared to an IdP that does not) simply due to a reduced attack surface as described in [37].
The fact that unsigned authorization requests do not result in serious drawbacks can be explained by the fact that two-way authentication is established during artifact resolution before the assertion is delivered.
Additionally, this use case provides assertion secrecy even though no assertion encryption mechanism is used. Again, this fact can be explained by the artifact resolution mechanism.
Hence, it is hard for SAML users to decide what optional security mechanisms to include since it is highly non-trivial to deduce the impact of those security mechanisms on the final SSO system.
VI Related work
Considering that SAML-based SSO systems are widely used, especially in government, corporate and educational settings, there are surprisingly few efforts to formally analyze these systems. Our research is the first attempt at a formal analysis of the POST/Artifact bindings use case, as well as the first attempt to analyze the technical specification rather than a particular deployment.
The most notable example of a formal analysis of a SAML SSO use case is the research of Armando et al. [18, 19, 20].
They model and analyze the Redirect/POST use case, as well as an actual deployment by Google, using a SAT-based model checker SATMC [39]. They consider two security properties — authentication based on non-injective agreement and the resource secrecy. The most significant contribution is the discovery of a crucial flaw in the Google’s deployment — the SP identifier was left out from the <AuthnRequest> message leading to an attack breaking both security properties considered. The authors also note that the tool fails to find similar attacks when the SP identifier is included in the message as required by the SAML specifications.
In the only other relevant attempt at a formal analysis of SAML use cases [40], Ferdous and Poet analyze the same use case as in [20], and create a simpler model, but one that includes an explicit user authentication step and a rudimentary digital identity model. The security properties are again similar to [20] and include basic authentication and secrecy properties. They use the High-Level Protocol Specification Language (HLPSL) to specify the protocols and the Automated Validation of Internet Security Protocols and Applications (AVISPA) tool with the On-the-fly Model-Checker (OFMC) backend for analysis. Besides the SAML use case, they also investigate other SSO protocols namely OpenID Connect and OAuth. The analysis by the tool discovered no attacks, but the authors do not make explicit claims on the exact setting in which the properties hold.
Our work differs from the two efforts mentioned above in several aspects. First, we focus on a different SAML use case. More importantly, we build a much more detailed and comprehensive model of the system (the protocol setup, the protocol variants, a more general adversary that can compromise any participant, etc.) and consider a larger number of relevant security properties. Finally, the bounded model checker used in [20] only guarantees the absence of attacks assuming a small fixed number of participants and sessions. In contrast, our positive results hold in a much more general setting with an unbounded number of protocol participants and sessions.
In the more general area of formal analysis of Web SSO systems and standards, Fett et al. develop an expressive Dolev-Yao style model of the web infrastructure (the FKS model) and use it to perform a comprehensive formal analysis of OAuth [41] and OpenID Connect [42]. The analysis is not machine-verified, but others have replicated their models and confirmed the results using the Tamarin prover [35, 38].
Finally, there are several important efforts to analyze particular implementations and deployments of SAML-based systems and detect low-level attacks. Somorovsky et al. analyze 14 major SAML frameworks and show that 11 of them have vulnerabilities related to XML signature handling [15]. Mainka et al. perform security analysis of 22 SaaS cloud providers that use a SAML-based SSO [16]. They identify eight different SAML-related attack vectors and show that 90% of the providers are vulnerable to at least one attack. Finally, Engelbertz et al. identify numerous vulnerabilities in SAML-based eID implementations in European Union countries [17]. They also develop the EsPReSSO tool for automatic security analysis of SAML-based SSO systems.
VII Conclusions and future Work
We deliver a formal model of the use case from the SAML V2.0 Web Browser SSO Profile: SP-initiated SSO with POST/Artifact Bindings. While building our model we used specific formal model of TLS communication which enables Dolev-Yao adversary to exchange messages with honest agents by compromising an honest agent. We analyzed 8 protocol variants that emerged from differently defining three sensitive optional features. We tested these variants against comprehensive set of authentication, secrecy and freshness properties. Automatic verification analysis has found the artifact-session confusion attack which breakes the IdP-SP non-injective agreement property. Although by this attack the adversary could gain access to arbitrary user resource, we argued that there is no practical impact of this attack. But this attack emphasises that care must be taken while implementing SAML processing rules completely.
In future work we would like to create comprehensive formal models of all possible use cases of Web Browser SAML SSO Profile and to automatically verify presented security properties. In such work we would like to include analysis of concrete SAML SSO implementations and to compare those implementations with our formal model.